OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = –60

OpenStudy (anonymous):

use the equation of  then you have -60=12 +(13-1)d simplify -60=12+12d then solve for d, d = -6 then you you have d, and after that use the equation to look for the 32nd term